3) Use simple fixed-point iteration to locate the root of f(x) = 2 sin(x) - x
Use an initial guess of Xo = 0.5 and iterate until Eg s 0.001%. Verify that the process is linearly convergent.
Answer #1

f(x) = 0 => x = 2*sin(\sqrt{x})

according to simple fixed-point iteration method:

In+1 = 2 sin( n)

we may create the following table in MS excel:

xn x n+1 e (xn+1 - xn) en+1/en 0.5 1.299274 0.799273878 - 1.299274 1.817148 0.517873626 0.64793 1.817148 1.950574 0.133426413

Hence, the converged root is: 1.97238 _________________ Answer

Since n+1 = 0.1179 = constant (close to convergence) En ,

hence, we may conclude that the convergence is linear in nature.